OpenStudy (highschoolmom2010):

write an equation for a circle with the given conditions area of 36 pi & a center at point (4,7)

OpenStudy (anonymous):

you can calculate radius r by \[\pi r^2= 36 \pi\] and after that use eqn of circle \[(x-a)^2 +(y-b)^2 =r^2\] where (a,b)=(4,7)

OpenStudy (jdoe0001):

as @gyanu said using the Area of a Circle, solve for "r" to get the radius, once you get the radius you're already given the center of the circle and that's all you need for the circle's equation, thus just plug them all in :) like so -> http://www.mathwarehouse.com/geometry/circle/images/equation-of-circle/example-2-equation-of-circle.png